1. a. The pressure inside a hydrogen-filled container was 2.10 atm at 21 ∘C. What would the pressure be if the container was heated to 87 ∘C ?

b. At standard temperature and pressure (0 ∘C and 1.00 atm ), 1.00 mol of an ideal gas occupies a volume of 22.4 L. What volume would the same amount of gas occupy at the same pressure and 95 ∘C ?

c. Calculate the number of molecules in a deep breath of air whose volume is 2.15L at body temperature, 36 ∘C, and a pressure of 735 torr.
Answer #1

O Intel Instant Find Firend presure of a gas - 2. 10lm , temporaline (71) = (21+ 273 K = 294k. tempoerture (g) = (87+273 ) =o volume (V) 2.15L Temparature (1) = (36+273) K = 3091 preane (o)= 735 torr = 735 an ( 760 tott alen] 0.9671 ahn from Ideal g